Rundle Company produces commercial gardening equipment. Since production is highly automated, the company allocates its overhead costs to product lines using activity-based costing. The costs and cost drivers associated with the four overhead activity cost pools follow: Activities Unit Level Batch Level Product Level Facility Level Cost $ 65,000 $ 18,900 $ 15,000 $ 270,000 Cost driver 2,600 labor hrs. 35 setups Percentage of use 15,000 units Production of 720 sets of cutting shears, one of the company’s 20 products, took 240 labor hours and 8 setups and consumed 19 percent of the product-sustaining activities. Required Had the company used labor hours as a companywide allocation base, how much overhead would it have allocated to the cutting shears? How much overhead is allocated to the cutting shears using activity-based costing? Compute the overhead cost per unit for cutting shears first using activity-based costing and then using direct labor hours for allocation if 720 units are produced. If direct product costs are $150 and the product is priced at 25 percent above cost for what price would the product sell under each allocation system?
Answer #1

a. Single overhead rate = 368900/2600 = $141.88 per labor hour

Overhead allocated = 240 x $141.88 = $34051
Overhead per unit = $34051/720 = $47.29 per unit

b.

Overhead Activity Cost Quantity of cost driver Activity Rate
a b c d=b/c
Unit level $                65,000.00 2600 labor hours $                        25.00 per labor hour
Batch level $                18,900.00 35 Number of Setups $                      540.00 per setup
Product level $                15,000.00
Facility level $             270,000.00 15000 units $                        18.00 per unit
$             368,900.00
Overhead allocation
Cutting Shears
Unit level $                  6,000.00 =240*25
Batch level $                  4,320.00 =8*540
Product level $                  2,850.00 =15000*19%
Facility level $                12,960.00 =720*18
Total $                26,130.00
Total Units 720
Overhead cost per unit $                        36.29

c.

ABC Labor hours
Allocated Overhead $                        36.29 $                        47.29
Direct Cost $                      150.00 $                      150.00
Total Cost per unit $                      186.29 $                      197.29
Desired Profit $                        46.57 $                        49.32
Selling price $                      232.86 $                      246.61